### v2

# ISP-AD: A Large-Scale Real-World Dataset for Advancing Industrial Anomaly Detection with Synthetic and Real Defects

Abstract
Automatic visual inspection using machine learning-based methods plays a key role in achieving zero-defect policies in industry. Research on anomaly detection approaches is constrained by the availability of datasets that represent complex defect appearances and imperfect imaging conditions, which are typical to industrial processes. Recent benchmarks indicate that most publicly available datasets are biased towards optimal imaging conditions, leading to an overestimation of the methods’ applicability to real-world industrial scenarios. To address this gap, we introduce the Industrial Screen Printing Anomaly Detection dataset (ISP-AD). It presents challenging small and weakly contrasted surface defects embedded within structured patterns exhibiting high permitted design variability. To the best of our knowledge, it is the largest publicly available industrial dataset to date, including both synthetic and real defects collected directly from the factory floor. In addition to the evaluation of defect detection performance of recent unsupervised anomaly detection methods, experiments on a mixed supervised training approach, incorporating both synthesized and real defects, were conducted. Even small amounts of injected real defects prove beneficial for model generalization. Furthermore, starting from training on purely synthetic defects, emerging real defective samples can be efficiently integrated into subsequent scalable training. Research findings indicate that supervision by means of both synthetic and accumulated real defects can complement each other, meeting demanded industrial inspection requirements such as low false positive rates and high recall. The presented unsupervised and supervised dataset splits are designed to emphasize research on unsupervised, self-supervised, and supervised approaches, enhancing their applicability to industrial settings.
keywords:
Industrial Anomaly Detection, Industrial Anomaly Detection Dataset, Surface Defect Detection, Synthetic Defects, Automatic Visual Inspection
1 Introduction
With the affordability of modern computing power, the research and subsequent application of deep learning-based surface defect detection in industry is on the rise [54, 31]. As part of smart manufacturing and emerging Industry 5.0, recent publications address surface defect detection as an anomaly detection problem [42, 44, 69]. In the context of surface defect inspection, an anomaly can be defined as any unwanted deviation from the sample’s permitted surface variability and appearance, which could have known and unknown characteristics. Furthermore, the rare occurrence of anomalies compared to fault-free instances leads to heavily imbalanced data distributions [3]. These anomalies can affect both the aesthetic and functional properties of the product and are therefore of major interest to industrial quality control. Up to now manual inspection by humans is still part of industrial quality control. This repetitive task is prone to subjective assessment and fatigue resulting in quality fluctuations [16, 37]. Integrating automated visual inspection processes into the factory floor enables improved product quality, efficiency by reducing human labor, their subjective assessment as well as subsequent production costs. Furthermore, its application is not limited to specific industries and is already used in the e.g. automotive, textile, electronics and agriculture industries [55].
A key element of any machine vision system is its underlying defect detection algorithm, enabling the decision-making process of categorizing a sample as fault-free (normal) or defective (anomalous) [56]. The performance of machine-learning and more specific deep learning-based methods is dependent on the quality and amount of available data. Supervised algorithms based on deep convolutional neural networks (DCNN) [38] have gained remarkable performance in various defect detection tasks, relying on large labeled datasets [78, 61].
However, collecting large amounts of fault-free and defective data and subsequent labeling is labor-intensive and often impractical in many industrial scenarios. Additionally, supervised methods struggle to generalize to unseen defects that were not part of the training data. To address these issues, current research focuses on unsupervised anomaly detection methods [18]. These methods solely rely on fault-free samples during training, learning feature representations of its underlying normal data distribution. With the publication of datasets such as MVTec [8], a wide range of different approaches have emerged, usually based on the comparison of image features or the reconstruction of normal image instances.
Extracted features, obtained by means of layers of e.g. DCNN, are applied in memory bank [39, 59, 70], normalizing flow [80, 72, 64] or knowledge distillation-based approaches [9, 60, 6]. During inference, extracted test features are compared to learned normal representations using distance metrics or distribution mappings.
Reconstruction-based methods using autoencoders [10] or generative adversarial networks (GAN) [77] attempt to reconstruct normal image regions while failing in resembling anomalies, resulting in anomaly scores.
These approaches suffer from poor reconstruction performance on fine-grained structures as well as demanding model training (mode collapse in GANs). Recently, Diffusion-based approaches [65, 48, 75] have gained increased attention, to address reconstruction limitations. These methods utilize iterative noising and denoising processes to model underlying data distributions at a computationally intensive cost.
Additional approaches are based on the synthetization of defects in both image and/or feature spaces, showing improved detection performance on image and pixel-level tasks [42, 73].
Due to the inherent nature of manufacturing processes, defective samples accumulate over time. Despite the promising detection performance of unsupervised methods, their ability to utilize defective data is limited. Consequently, incorporating both real and synthesized defects into the training process is emerging as a viable strategy, enhancing the method’s discriminative capabilities.
In addition to defect detection performance, the industrial applicability of a defect detection method depends on meeting process requirements such as process cycle times, adaptability to different products, and robustness to permitted design variability and varying operational conditions. To evaluate model performance in an industrial setting, a dataset that reflects these conditions is crucial. However, a review of the literature shows that most publicly available datasets are generated under “laboratory conditions”, failing to capture the complexities of industrial environments [42, 2]. As a result, the benchmarks of state-of-the-art (SOTA) anomaly detection methods on these datasets are often overestimated compared to real-world industrial scenarios.
1.1 Contributions
The aim of this publication is twofold. First, to bridge the gap to industry, we introduce a dataset derived from a real-world industrial use case in screen printing. This dataset provides both unsupervised and supervised training data, comprising synthetic and real defects collected during production. Second, we investigate a mixed supervised training approach that efficiently utilizes available weakly labeled data consisting of both, synthetic and real defects. By means of this publicly available data set, further research in leveraging synthetic as well as real defective data for industrial anomaly detection should be emphasized.
Thus, the main contributions can be summarized to:
  • Introduction of a novel dataset of structured patterns, captured using three different optical modalities. The dataset originates from a real-world industrial use case in screen printing, including permitted process-specific data variability. With a total of 312 674312674312\,674312 674 fault-free samples and 246 375246375246\,375246 375 defective samples, where 245 664245664245\,664245 664 are synthetic and 711 are real, it represents the largest industrial defect detection dataset to date, enabling both unsupervised and supervised training scenarios.
  • Image and pixel-level benchmarks of SOTA unsupervised approaches on the proposed demanding imaging modalities, containing small and weakly contrasted defects with high permitted design variability.
  • The mixed supervised training procedure introduced by [26] was further investigated by utilizing various proportions of injected real defects. Starting from purely synthetic training, the incremental incorporation of accumulated weakly labeled real defects led to a notable improvement in defect detection performance, particularly for challenging small and weakly contrasted defects. Additionally, the limited use of previously unseen defects significantly enhanced generalization performance, complementing synthetic training data. Thus, emerging defective samples with previously unseen feature distributions can be efficiently integrated into the mixed training procedure.
  • Novel industrial benchmark, enabling research on unsupervised, self-supervised and supervised anomaly detection approaches.
2 State of the Art
2.1 Supervision in Industrial Anomaly Detection
Unsupervised methods do not rely on defective samples during training, avoiding the risk of bias towards seen anomalies that can occur in supervised settings. However, the absence of knowledge about anomalous data results in a lack of discriminative features, making it challenging to distinguish subtle anomalies from normal data [71, 21]. This can result in false positives for normal samples with high permitted variability or overlooked defective areas.
To address this, augmentation methods have emerged that synthesize defective samples, enabling models to learn more discriminative features by incorporating these synthetic representations during self-supervised training tasks. Methods like CutPaste or NSA [63, 40] synthesize defective samples by cropping patches from normal samples, augmenting them (e.g., resizing or rotating), and either pasting them or blending them into random positions using techniques like Poisson image editing [49].
In contrast to sampling from the examined data distribution, DRAEM [74] generates diverse defect shapes by extracting textures from different domains (out-of-distribution) [17] using masks generated via Perlin noise [50]. Additional approaches use gaussian or simplex noise [51] added onto normal images to generate synthetic defects [67, 13]. More realistic defects, referred to as in-distribution defects, can be achieved by generating defect textures on normal samples by means of random walks [25].
Recent trends leverage generative models such as GANs and diffusion-based approaches for defect synthetization [24, 79, 23, 28, 30] GAN-based approaches rely on sufficient training data, including defective samples, and often struggle to synthesize fine grained patterns.
In contrast, denoising diffusion models aim to address these issues by creating more realistic defects while requiring fewer defective samples.
Besides defect synthetization in the image domain, recent methods have shown promising results in generating defects directly in the feature space, e.g., by introducing gaussian noise [43] or sampling from a set of codebook features [73]. Global and Local Anomaly co-Synthesis Strategy (GLASS) [14] combines constrained defect synthetization at both feature and image levels, resulting in "near-in" and "far-from" normal sample distribution anomalies, achieving SOTA performance on MVTec [8].
However, common problems of image-level synthetization are the lack of realism and diversity, while feature-level synthetization is hard to control but more efficient. As a consequence, additionally leveraging real defective samples accumulated during production is a complementary approach to increase feature diversity.
Research using both synthetic and real samples during training has been reported in binary classification and object detection [53, 52, 26, 20] as well as in recent supervised anomaly detection approaches [57, 76, 21].
For instance, [53] generated synthetic defects using a GAN and trained an object detection architecture with both real and synthetic data, achieving a 5 % improvement in average precision compared to using only real defects. Similarly, [20] reported enhanced defect detection performance on a small-sized dataset by additionally incorporating synthesized defects generated via a denoising diffusion approach. FuseDecode [34] proposes a novel autoencoder-based anomaly detection model initially trained on noisy, unlabeled data. The predictions assist in generating weakly labeled datasets, enabling mixed supervision with synthesized and collected real defects, thereby reducing labeling efforts.
Compared to supervised classification approaches, supervised anomaly detection methods reduce bias toward seen real defects by incorporating strategies such as out-of-distribution defect synthesis.
2.2 Industrial Anomaly Detection Datasets
In addition to efforts in dataset preparation steps, including image capturing, data cleaning, preprocessing, and labeling, obstacles such as compliance standards for the investigated product impede dataset publication. As a result, many publicly available datasets are generated in laboratory settings, attempting to mimic industrial use cases.
Furthermore, due to the rare occurrence of defective samples, the artificial generation of defects is emphasized. This synthetization can be conducted at the image-level using algorithms (e.g., generative models or 3D rendering [19]) or manually on physical samples using appropriate tools.
To mitigate the need for defective samples during training, many industrial datasets are designed for unsupervised settings. In such a setting, defective data is only included in validation and/or test splits, minimizing efforts in defective data preparation.
Table 1: Comparison of SOTA anomaly detection datasets by sample size, defect data availability, and generation method (manual, synthetic, real-world). The proposed dataset, ISP-AD, introduces supervised and unsupervised data splits with synthetic and real defects, collected at the factory floor.
Dataset #Total #Good #Defects #Defects Train #Classes Modality Defect Generation
MVTec $5354$ $4096$ $1258$ 0 15 rgb/mono manually
VisA 10 8211082110\,82110 821 $9621$ $1200$ 0 12 rgb manually
Real-IAD 151 050151050151\,050151 050 99 7219972199\,72199 721 51 3295132951\,32951 329 0 30 rgb manually
RAD $1510$ $286$ 1224 0 4 rgb manually
DAGM 16 1001610016\,10016 100 14 0001400014\,00014 000 2100 1050 10 mono synthetic
MIAD 105 000105000105\,000105 000 87 5008750087\,50087 500 17 5001750017\,50017 500 0 7 rgb synthetic
KolektorSDD2 $3335$ $2979$ 356 246 1 rgb real-world
Textile $2150$ $1814$ 336 0 6 rgb real-world
BTAD $2830$ $2250$ 580 0 3 rgb real-world
VAD $5000$ $3000$ 2000 $0001$ rgb real-world
ISP-AD 559 049559049559\,049559 049 312 674312674312\,674312 674 711/245 664245664245\,664245 664 446 3 rgb/mono real-world/synthetic
Table 1 visualizes common industrial anomaly detection datasets. The MVTec dataset [8] has significantly impacted the field by introducing an unsupervised anomaly detection dataset consisting of 15 industrial categories, grouped into 5 texture and 10 object-based classes. The dataset contains a total of 5354 images, including 1258 defective samples with ground truth masks that enable pixel-level evaluation. Images were mostly captured under highly controlled illumination conditions, and defects were manually generated to produce realistic visual appearances.
The VisA dataset [81] extended these efforts by including objects with complex structures, such as printed circuit boards, and multiple instances resulting in a dataset twice the size of MVTec. Considering additional industrial scenarios, Real-IAD [68] introduced a multi-view dataset, covering 30 classes of a variety of materials, such as plastic, wood, ceramics etc. In addition, RAD [15], included uneven illuminations and blurry collections to imitate varying real-world inspection conditions.
Despite advancements in dataset size and increased data variability, the data generation process (including defective samples) was still performed manually, leaving domain gaps compared to factory-floor conditions.
Both DAGM [46] and MIAD [5] are based on synthetic defect generation. DAGM, published in 2007, contains 10 synthetically generated classes summing up to 14 0001400014\,00014 000 fault-free background textures and 2100 defectives, one for each background texture. In contrast, the more recent MIAD contains 105 000105000105\,000105 000 images of various outdoor industrial maintenance scenarios. MIAD leverages 3D graphics software [19] to render realistic 3D scenes with varying surface textures, backgrounds and viewpoints on both fault-free and defective objects.
In contrast to above described industrial datasets, including artificial generated defects, samples collected at industrial production lines are of particular interest. The KolektorSDD2 [12] dataset addresses a practical real word example, containing complex textured background structures. Defects in this dataset vary in size and shape, ranging from small scratches to large surface imperfections.
Another example is a textile dataset [66], further referred to as Textile, that reflects imperfect industrial conditions, such as image blurring and environmental contamination. A further inspection application of three different industrial products, showcasing body and surface defects, is presented within the BTAD dataset [47].
The recently published VAD dataset [4] bridges the gap to supervised anomaly detection by additionally introducing 1000 defective training images, thereby extending the unsupervised setting. The investigated piezoelectric element in VAD is prone to structural defects (e.g., cracks or pollutions) as well as to logical defects (e.g. wire or solder position).
3 Industrial Screen Printing Anomaly Detection Dataset: ISP-AD
Benchmarks on anomaly detection datasets such as MVTec [8] already show high detection performance above 99 % image-level Area Under the Receiver Operating Characteristic curve (AUROC), indicating less demanding inspection scenarios for SOTA anomaly detection methods.
Therefore, recent efforts in dataset generation focus on mimicking real-world inspection conditions by altering viewpoints, illumination, background, or product placement. However, a review of the literature still reveals a gap in the availability of large-scale datasets (several 10 thousand instances) captured at real-world production lines. Compared to clearly pronounced defect classes in certain object categories [8, 47, 81], introducing small-scale and subtle defects is important for creating more challenging datasets. Additionally, most datasets are prepared for unsupervised settings, not considering additional defective training samples for supervised approaches.
The proposed industrial dataset aims to address known limitations by introducing the following features:
  1. [1.]
  2. 1.
    Data captured from a real-world industrial manufacturing scenario.
  3. 2.
    Small and weakly contrasted defects on structured patterns, relative to large sample areas.
  4. 3.
    Three imaging modalities to enhance data variety.
  5. 4.
    Large-scale industrial data splits applicable to both unsupervised and supervised settings. Test splits are imbalanced, typical to manufacturing data distributions (zero-defect policies).
  6. 5.
    Synthetic and collected real defective data accumulated during production, suitable for additional supervision.
  7. 6.
    High permitted sample variance arising from the sample itself, preprocessing, and imperfect imaging conditions.
This section provides a brief overview of the sample under investigation and its defect classes. Additionally, the process of dataset generation, its specifications, and limitations.
3.1 Industrial Screen Printing and Defect Classes
[…]
Figure 1: Examples of punctual defects from the "Points" defect group, their pixel-level ground truth masks (GT), and overlays (Points + GT). Each row corresponds to a certain defect (fiber, pinhole, mechanical deformation) within its imaging modality (top to bottom: LSM-1, LSM-2, ASM). The pinhole defect in row 2 appears small (few pixels in extension), whereas defects like the fiber in row 1 exhibit low contrast within high sample variability.
Figure 2: Examples of defects from the "Area" defect group, their pixel-level ground truth masks (GT), and overlays (Area + GT). Each row corresponds to a defect (grid, pattern misalignment, squeegee stroke) within its imaging modality (top to bottom: LSM-1, LSM-2, ASM). The pattern misalignment and grid defects exhibit distinct features across the entire patch, while the squeegee stroke appears within the transition area of direct reflection (bottom row).
3.2 Dataset Generation
The generation of appropriate datasets builds the basis for achieving the required defect detection performance of machine learning based methods applied on the production line. Therefore, the design of validation and test splits mimicking the inspection process with its imbalanced data distributions is of utmost importance to perform reliable evaluations. Furthermore, generated data distributions must resemble the imperfect imaging conditions caused by environmental influences (e.g., varying illumination), data processing (varying background, viewpoint and segmentation artifacts), as well as permitted sample variance.
The proposed dataset is based on the work presented in [35], including its efficient data preprocessing and labeling workflow. Prior to image acquisition, fault-free and defective samples were pre-sorted by domain experts. The sample state is characterized as fixed, thus additional temporal sample alterations can be excluded. Moreover, the samples were selected from different production batches, resembling the permitted design variability. Images were acquired by means of all three optical modalities (LSM-1, LSM-2, ASM) of the inspection system demonstrator installed on the factory floor. The acquired images of fault-free and defective samples were assigned to separate training and test splits. This step avoids any unwanted correlation of fault-free and defective patches during the subsequent patch extraction process.
The extraction of augmented fault-free training patches (256 x 256 px) at random positions within the region of interest (ROI) is integrated into an automated procedure (steps 1 to 5 in section 5.1 [35]). Furthermore, following the central cropping in step 4, synthetic defects according to section 5.2 are generated in 50 % of the fault-free patches. The utilized algorithm [25] is based on a stochastic process resembling a random walk with momentum. Thus, the randomly generated defect textures enabled the synthetization of a wide range of punctual defect appearances, in all three optical modalities (Fig. 3). As a result, balanced supervised training datasets were obtained.
In addition, real defective patches were manually extracted on selected defective samples in the training and test splits. Thereby, central defect labeling with patch sizes of 512 x 512 px ensures proper defect positioning within augmented patches, which are cropped to a final size of 256 × 256 px during training.
Above described workflow minimizes the manual labeling effort to known defective samples and regions. The underlying image-level labels can be considered ’weak’ since no ground truth masks are provided, thereby eliminating the need for elaborate pixel-level annotation. In general, patch-based processing is an efficient augmentation technique in upscaling available data, enabling extraction of several 100 000100000100\,000100 000 patches from a few acquired samples with large FOVs.
Figure 3: Synthesized defects generated using the algorithm proposed in [25]. Hyperparameters were adjusted to simulate real punctual defects (e.g., pinholes in the middle) and elongated defects (e.g., bright contrasted fiber on the left; small mechanical deformation on the right patch) across all imaging modalities (left to right: LSM-1, LSM-2, ASM).
Fault-free test data was generated using a sliding window with a stride of 160 px, extracting overlapping patches (256 x 256 px) to cover a grid. The same patch extraction process was applied to the training data for the unsupervised methods.
As an additional data cleaning step, unwanted defective patches (e.g., contamination in the process environment or defects overlooked by domain experts) were removed from the unsupervised training and general test data.
As described earlier, the automatically extracted fault-free supervised training data was not further screened and can therefore be considered noisy. Based on a manual inspection of 500 randomly extracted and preprocessed (contrast enhanced) patches from LSM-1, it was estimated that 2 - 5 % contained overlooked defects, ranging from borderline cases to clearly contrasted defects such as dust and fibers.
To enable the evaluation of detection performance at the pixel-level, ground truth masks for the test patches were generated using the ilastik labeling tool [7], which is based on a random forest classifier. Following the assignment of pixel object classes, additional image processing techniques (e.g., erosion and dilation) were applied to ensure distinct defective regions. Examples of selected defective patches, ground truth masks (GT), and their overlays are shown in Fig. 1 and 2.
3.2.1 Data Splits
The aim of this publication is to advance research in solving real-world industrial inspection problems by leveraging both synthetic and available real defects accumulated during production processes. Therefore, supervised and unsupervised datasets for each optical modality, acquired from the factory floor, have been created. The supervised training data splits include up to approximately 100 000100000100\,000100 000 augmented fault-free patches and 100 000100000100\,000100 000 synthetic defective patches. Additionally, real defective patches, extracted from defective samples, are categorized into two distinct defect groups: area and points. For unsupervised approaches, the fault-free training data includes up to 3678 patches in LSM-1, with reduced splits containing a maximum of 500 patches available for all modalities. These smaller splits are designed to minimize computational effort for e.g. memory demanding methods. The test data for both supervised and unsupervised approaches is identical, mimicking typical imbalanced inspection data distributions. Table 7 summarizes the proposed dataset with its synthetic and real defects available for supervision.
3.2.2 Data Format
In general, LSM-1 and LSM-2 patches are provided as 8-bit (rgb) .png files, while ASM patches are provided as single-channel 8-bit .png files. Ground truth masks are stored as single-channel 8-bit .png files. In the case of supervised training data, all patches are stored as single-channel 8-bit within the .hdf5 large file storage format. In addition to the image-level labels, pixel-level GTs are available due to defect synthetization. The folder structure of the unsupervised datasets follows the official MVTec [8] scheme.
3.2.3 Augmentation Settings
According to step 3 in section 5.1 [35], the following augmentation settings were applied to the supervised training data: vertical and horizontal flipping, rotation, shearing, and scaling. Random translation is not included in this list, as it is part of the prior random patch extraction process. Random affine transformations were applied to 80 - 95 % of the extracted patches. In addition, random contrast and brightness adjustments were applied. The corresponding transformation parameters with its values and ranges are shown in Table 8.
3.2.4 Additional Data
In addition to proposed data splits, supplementary fault-free patches for all modalities are provided. These patches are the unaugmented counterparts to the fault-free supervised training data described above. They consist of 20 0002000020\,00020 000 and 10 0001000010\,00010 000 fault-free patches in rgb format for LSM-1 and LSM-2, and 20 0002000020\,00020 000 patches in single-channel format for ASM, all stored as .png files.
3.2.5 Limitations
Due to the limited amount of defective data available, a validation split was omitted. Although the fault-free patches in the test splits were extensively screened, it cannot be guaranteed that the labels are entirely defect-free.
In ASM, the manual extraction of defective patches introduced additional masked borders on the left and right patch sides (a few pixels in width), which could potentially lead to false positive predictions in non-robust unsupervised approaches, such as reconstruction-based methods. Therefore, cropping these regions during evaluation is recommended to mitigate this effect. Furthermore, the same defects may appear in different patch positions. However, due to varying illumination conditions in this modality, this can be considered an additional form of augmentation.
4 Defect Detecion Methods
The recently published work [35] introduced an efficient data preprocessing workflow utilizing weakly labeled defective data in a supervised training approach. Furthermore, the possible extension to a mixed training approach using both synthetic and available real defective data [26] was outlined, enhancing defect detection sensitivity of overlooked defect classes.
Therefore, the evaluation of the mixed training approach on the presented demanding imaging modalities should give further insights in leveraging supervised methods in industrial anomaly detection. For comparison several SOTA unsupervised anomaly detection methods are investigated. Utilized approaches as well as their training settings will be described in the following section.
4.1 Mixed Supervised Training
In industrial production, defects emerge at various stages of the manufacturing process and may be unavailable during training or even previously unknown. Thus, defects can be categorized as:
  • Seen defects: Known defect classes that are available during training (synthetic or real).
  • Unseen defects: Unknown or unavailable defect classes that may exhibit previously unobserved feature distributions.
The proposed mixed supervised training approach showed increased detection performance by leveraging both synthetic and real defects during training, compared to training separately on synthetic or real ones [26]. Thus, depending on the availability of real defects, the method can be applied in cold-start scenarios, starting with purely synthetic training and transitioning to a mixed setup as real defects accumulate. This is an important capability in industrial applications, as the model can be improved step by step with newly available defective samples.
However, the ability to synthesize certain defect classes using the algorithm in [25], such as squeegee strokes shown in Fig. 2, is limited or even infeasible. Thus, even small amounts of previously unseen real defects may contribute to learning feature representations that could not be captured by synthetization or were absent in the available real data. Additionally, incorporating real defects, similar to synthesized ones, increases the feature diversity of already seen defects.
Utilizing the above introduced datasets a DCNN, pretrained on ImageNet [32], is trained in a supervised manner. Synthetic defective patches were generated in 50 % of the fault-free patches. During training, each patch in this balanced stream is replaced by an augmented real defective patch with a probability of considerably less than 50 %. Thus, the objective is to complement the synthetic defects, not to replace them.
Supervised binary classification is inherently a closed-set problem, in contrast to the open-set problem of anomaly detection described above. Adapting the classifier to emerging, previously unseen defect classes with a minimal set of recently available samples would be an important capability in industrial settings.
The following experiments in section 5.2 aim to investigate the generalization capabilies of the proposed mixed training approach for previously unseen defects (e.g. that were not able to be synthesized). In addition, its scalability is examined under varying proportions of collected real defects within all imaging modalities: LSM-1, LSM-2 and ASM.
4.2 Training Settings
[…]
4.3 Unsupervised Methods
In order to compare the defect detection performance to SOTA unsupervised approaches, methods for feature embedding, reconstruction and synthetization are investigated. Thus, the methods used are grouped according to their approaches and thus briefly described. In addition, method and training settings utilized in the experiments are listed.
4.3.1 Normalizing Flows
In contrast to utilizing large memory banks for feature comparison, normalizing flows model normal feature distributions. During training, the initial complex normal feature distribution is transformed into a standard normal distribution through a series of invertible mappings. In contrast to DCNNs as feature extractors, U-Flow [64] utilizes a multiscale vision transformer architecture [22], pretrained independently for each scale. A fully invertible architecture is achieved by adapting a UNet-like structure [58] to normalizing flows. Thus, the feature extractor acts as the encoder, while the normalizing flow serves as the decoder.
4.3.2 Student-Teacher
Within this approach, a shallow student network attempts to mimic the output of a pre-trained (distilled) teacher using knowledge distillation. Exclusively trained on fault-free data, the distilled student network fails to predict the teacher’s output on anomalous features, resulting in an anomaly score. Efficient-AD [6] provides an lightweight network architecture with restricted receptive field to accelerate feature extraction. Additionally, a hard feature loss function and a loss penalty are introduced to prevent the student from generalizing its imitation to out-of-distribution images. Furthermore an integrated autoencoder ensures detection of logical anomalies.
4.3.3 Denoising Diffusion
Denoising diffusion probabilistic models are a class of generative models inspired by non-equilibrium thermodynamics. They consist of a forward diffusion process that gradually adds gaussian noise to the input image over several steps, transforming it into a standard normal distribution. During training, the reverse diffusion process learns to denoise the corrupted image step by step by minimizing the difference (mean squared error) between the predicted noise and the actual noise added during the forward process. The method proposed in Denoising Diffusion Anomaly Detection (DDAD) [48] utilizes target images to guide the denoising process, improving its ability to reconstruct normal patterns. Anomaly localization is achieved through both feature-wise and pixel-wise comparisons between the reconstructed image and the input image. The pretrained feature extractor is adapted to the target domain using examples generated by the denoising model, further enhancing detection performance.
4.3.4 Defect Synthetization
GLASS [14] extends purely image-level anomaly synthetization by introducing Global Anomaly Synthesis (GAS) at the feature level and Local Anomaly Synthesis (LAS) at the image level. GAS utilizes gaussian noise guided by gradient ascent and truncated projection to synthesize anomalies in the feature space near the normal sample distribution (in-distribution anomalies). LAS generates out-of-distribution anomalies by overlaying textures on normal images using Perlin masks, similar to the approach in [74]. Finally, a segmentation network is trained end-to-end using three loss functions: normal feature, local anomaly feature, and global anomaly feature loss. During inference, only the normal feature branch is utilized for anomaly detection.
4.3.5 Implementation Details
To standardize the evaluation workflow, the Anomalib API [1] was utilized for the following unsupervised methods [64, 6]. Therefore, a training and evaluation pipeline, was implemented. Default method settings, as documented in the model configuration files, were applied. In case of Efficient-AD, the medium (M) patch descriptor network was chosen.
Additionally, the official implementations of DDAD [48] and GLASS [14], along with their configuration files, were selected. The target image conditioning parameters $w$ and $w\_DA$ were set to 3, together with 4 domain adaptation epochs, as specified in DDAD. In GLASS, the manifold hypothesis was selected as the GAS strategy, while foreground masks were omitted in LAS.
In general, the training input shape for image patches was set to 256 x 256 px. Despite center cropping (224 x 224 px) in DDAD for LSM-1, and center cropping (224 x 224 px) in ASM, no additional preprocessing was applied to the raw image patches. Due to Efficient-AD’s autoencoder architecture, the input shape was kept constant at 256 x 256 px. To enable feature extraction by means of the multi-scale vision transformer backbone in U-Flow, patches were resized to 448 x 448 px. Configuration files for all methods are publicly accessible in the dataset repository (Future Applications).
4.3.6 Training Settings
[…]
4.3.7 Hardware Setup
Experiments were performed on desktop workstations, equipped with NVIDIA GPUs (GeForce RTX 3090 or GeForce RTX 4080), AMD 16-core processors and 64GB RAM running on OS Windows 10/11. The Python environments (version $\geq$ 3.9.15) utilized GPU versions of the PyTorch framework (version $\geq$ 2.1.0) with CUDA Toolkit version 11.8, as well as PyTorch Lightning (2.4.0).
5 Experiments and Results
The following section investigates the defect detection performance of the supervised and unsupervised approaches proposed in section 4, applied across all optical modalities: LSM-1, LSM-2, and ASM.
Table 2: Defect detection performance of the investigated mixed supervised training approach [26] evaluated on LSM-1. Starting from purely synthetic training, real defective patches were progressively injected with a probability of 1/32 at varying fractions (1/16, 1/8, 1/4, 1/2, 1) of collected defects. The best-performing fraction, based on image-level MCC, is highlighted in bold.
Defect Group #Real Defects MCC Recall (%) FPR (%) TN TP FN FP
- 0 0.81 76.8 0.7 1459 73 22 11
points 10 0.86 82.1 0.5 1463 78 17 7
area 10 0.90 84.2 0.2 1467 80 15 3
mixed 10 0.91 84.2 0.1 1469 80 15 1
mixed 21 0.91 86.3 0.2 1467 82 13 3
mixed 43 0.93 91.6 0.3 1466 87 8 4
mixed 86 0.93 90.5 0.2 1467 86 9 3
mixed 172 0.96 95.8 0.2 1467 91 4 3
Table 3: Defect detection performance of the investigated mixed supervised training approach [26] evaluated on ASM. Starting from purely synthetic training, real defective patches were progressively injected with a probability of 1/32 at varying fractions (1/16, 1/8, 1/4, 1/2, 1) of collected defects. The best-performing fraction, based on image-level MCC, is highlighted in bold.
Defect Group #Real Defects MCC Recall (%) FPR (%) TN TP FN FP
- 0 0.70 62.2 0.6 1905 46 28 11
points 9 0.76 66.2 0.3 1910 49 25 6
area 9 0.91 91.9 0.4 1909 68 6 7
mixed 9 0.94 93.2 0.2 1912 69 5 4
mixed 18 0.92 91.9 0.3 1911 68 6 5
mixed 36 0.97 95.9 0.1 1914 71 3 2
mixed 73 0.96 94.6 0.1 1914 70 4 2
mixed 146 0.96 97.3 0.2 1912 72 2 4
5.1 Defect Detection Performance Metrics
In addition to commonly used metrics such as recall, false positive rate (FPR), and AUROC, metrics designed to handle class imbalance effectively were selected. For image-level metrics, the Matthews Correlation Coefficient (MCC) was applied. The MCC, which ranges from -1 (indicating inverse prediction) to 1 (indicating perfect prediction), incorporates all entries of the confusion matrix: true negatives (TN), true positives (TP), false negatives (FN), and false positives (FP). In this context, TN represents correct fault-free predictions, while TP represents correctly predicted defective patches. To account for pixel-level class imbalance, the per region overlap score (PRO) was chosen to assign appropriate weightings to different sized defective regions.
5.2 Mixed Supervised Training
To evaluate the scalability of the introduced mixed supervised training, different fractions of real defects accumulated during production were injected into the balanced stream of fault-free and synthetically generated defects.
Following random shuffling of all available defects, fractions of 1/2, 1/4, 1/8, and 1/16 were extracted in an ordered manner. This approach ensures that defects included in smaller fractions are also represented in larger ones. These fractions consist of both defect groups, points (Fig. 1 ) and area (Fig. 2), representing features that are either punctual or distributed across large sample areas. Additionally, to examine adaptability to unseen features, such as those missed during synthetization, training was performed exclusively on each defect group. In all experiments, starting from purely synthetic training, real defects were injected with a probability of 1/32.
To ensure comparability between experiments, the same training settings, including learning rates, were applied within each optical modality. Tables 2 and 3 visualize the image-level defect detection performance in LSM-1 and ASM.
5.2.1 Results
Training with purely synthetic defects in LSM-1 resulted in an MCC of 0.81 with an FPR of 0.7 %. As described in the previous publication [35], depending on the underlying preprocessing strategy, FPRs less than 0.59 resp 0.33 % were necessary to avoid unwanted false alarms, leading to unnecessary sample rejections.
Gradually adding real defective patches showed an increase in performance, achieving an MCC of 0.96 when utilizing all 172 available defects. This reduced the number of FN predictions from 22 to 4, representing a reduction of 81 %. By leveraging only 10 defective patches from both defect groups (mixed), an MCC of 0.91 was achieved. For comparison, supervised training with oversampling using all available (172) real defects, as described in [35], reached a comparable performance.
Training with separated fractions of point and area defects also showed an increase in detection performance, with area defects having more impact. Adding this defect group reduced group-related FNs from 6 in purely synthetic training to 1 in both mixed training and area-specific fractions. In mixed fractions greater than 1/8, no area defect was overlooked.
Training on the LSM-2 modality achieved near-perfect detection performance, overlooking only one defective patch. The injection of real defective patches resulted in faster convergence, requiring only 10 epochs compared to approximately 20 epochs for purely synthetic training.
The ASM modality can be defined as the most demanding dataset due to its specific illumination characteristics (see Fig. 2) and imperfect segmentation masks. Training on purely synthetic defects resulted in an MCC of 0.70, which could be significantly improved to greater than 0.95 by utilizing fractions upwards of 1/8. This improvement reduced FNs by up to 93 % through the injection of real defects. The FPRs were comparable to those observed in LSM-1, indicating the method’s robustness to permitted sample variations and challenging inspection conditions.
Leveraging the area defect group or mixed fractions had a significantly greater impact on improving detection performance compared to relying solely on punctual defects. In purely synthetic training, 23 out of 28 FNs were area defects. Training with injected area defects led to only a few overlooked (up to 3). Utilizing all 146 available defects resulted in just 2 FNs, one of each defect group.
As described earlier, the capability to synthesize area defects is limited. In the case of ASM, defect synthetization for the squeegee stroke defect class (Fig. 2) was not possible and can be characterized by having unseen feature distributions. In contrast, LSM-1 contains grid defects (Fig. 2) that exhibit similar punctual features to the synthesized ones. As already observed in [35], LSM-1 demonstrate greater generalization capability using synthetic defects, though ASM benefits more from the injection of area defects. For the ASM modality, the injection of 9 area defects led to an increase in the MCC by 0.21, reaching 0.91. The injection of similar synthesized feature distributions, as in the case of point defects, proves beneficial for the overall defect detection performance, including previously overlooked punctual defects.
Furthermore, training on several tens of thousands of fault-free patches introduced permitted normal feature variability, resulting in industry-applicable FPRs of less than 0.59 resp 0.33 %.
5.2.2 Limitations
The above results are based on the initial random selection of defective samples and are therefore expected to vary with different seeds during dataset generation. Extensive hyperparameter tuning and statistical analysis are left for future research. In addition, the distribution of available area and points defect groups is imbalanced, due to reasons of availablility (Table 7).
5.3 Unsupervised Methods
To estimate the defect detection performance of unsupervised methods in comparison to the supervised approach described above, evaluations were conducted using the same imbalanced test datasets. In addition to image-level labels, ground truth masks were utilized to assess the segmentation performance of each method. Performance metrics of various SOTA unsupervised approaches, applied to modalities: LSM-1, LSM-2 and ASM, are summarized in Tables 4 to 6.
5.3.1 Results
Table 4: Defect detection performance of investigated unsupervised approaches (Efficient-AD [6], U-Flow [64], GLASS [14], and DDAD [48]) on LSM-1. The best-performing method, based on image-level MCC, is highlighted in bold.
Method AUROC (%) MCC Recall (%) Precision (%) FPR (%) AUROC (%) PRO (%)
image-level pixel-level
Efficient-AD 92.6 0.87 82.1 94.0 0.3 80.8 4.1
U-Flow 95.9 0.77 75.8 80.9 1.2 99.1 28.3
GLASS 99.5 0.89 92.6 86.3 1.0 99.2 13.0
DDAD 91.9 0.60 63.2 62.5 2.4 93.7 16.8
Table 5: Defect detection performance of investigated unsupervised approaches (Efficient-AD [6], U-Flow [64], GLASS [14], and DDAD [48]) on LSM-2. The best-performing method, based on image-level MCC, is highlighted in bold.
Method AUROC (%) MCC Recall (%) Precision (%) FPR (%) AUROC (%) PRO (%)
image-level pixel-level
Efficient-AD 99.1 0.96 94.8 98.9 0.3 80.9 58.8
U-Flow 99.9 0.96 95.8 97.9 0.5 99.8 95.6
GLASS 99.5 0.94 96.9 93 1.8 99.9 99.2
DDAD 98.2 0.82 84.4 87.1 3.1 99.7 77.9
Table 6: Defect detection performance of investigated unsupervised approaches (Efficient-AD [6], U-Flow [64], GLASS [14], and DDAD [48]) on ASM. The best-performing method, based on image-level MCC, is highlighted in bold.
Method AUROC (%) MCC Recall (%) Precision (%) FPR (%) AUROC (%) PRO (%)
image-level pixel-level
Efficient-AD 85.5 0.31 28.4 38.9 1.7 88.4 4.5
U-Flow 92.4 0.49 52.7 48.8 2.1 89.3 19.9
GLASS 87.4 0.34 48.6 28.6 4.7 82.6 4.0
DDAD 90.9 0.38 55.4 29.7 5.1 96.6 6.5
Image-level defect detection performance in LSM-1 across different methods shows substantial variation with MCCs ranging from 0.60 at an FPR of 2.4 % to 0.89 at an FPR of 1.0 %. At the pixel level, minimum PRO-scores ranged from 4.1 to 28.3 %. In general, the low PRO-scores highlight the challenging task of detecting small punctual anomalies such as grid defects in Fig. 2. Common FPs occur at masked border regions or within areas of permitted variance in patterned structures (Fig. 4).
Figure 4: Illustration of the pixel-level detection performance (anomaly map overlay) for the evaluated SOTA unsupervised methods (Efficient-AD [6], U-Flow [64], GLASS [14], and DDAD [48]) on LSM-1, LSM-2, and ASM (top to bottom). The first column displays defects with their ground truth overlays (grid, pinhole, squeegee stroke). Overestimated defect areas (e.g., grid defect) and missed detections of small, low-contrast pinholes demonstrate the challenges of proposed industrial scenario. False positives frequently occur in masked border regions and areas of high contrast variations, as observed in ASM (bottom row).
The comparison of high pixel-level AUROCs ($>$ 99 %) compared to the measured PRO-scores illustrates their misleading interpretability in imbalanced scenarios.
Efficient-AD demonstrated a 7-fold reduction of FPR, compared to the reconstruction based approach (DDAD) based on denoising diffusion. Thus, Efficient-AD demonstrated an industry applicable FPR of 0.3 %, with qualitatively precise anomaly segmentations of small pronounced defects. However, despite the patch descriptor’s well-defined receptive field ensuring robustness, small defects remain boundary cases and are prone to misclassification. This is likely due to the bias introduced by ImageNet pretraining and the shallow patch descriptor architecture, which may limit its ability to extract fine-grained, in-domain features for subtle anomalies. In general, small and weakly contrasted point defects remain challenging even in GLASS, resulting in an recall of 92.6 %.
Compared to modality LSM-1, defect classes such as pinholes and pattern misalignments, which stand out prominently from the background pattern, exhibit more pronounced defect features. Consequently, both image- and pixel-level performance of various unsupervised approaches show improvements, achieving MCCs of up to 0.96 and PRO scores above 99 %.
However, segmentation performance varies significantly across the underlying approaches. Small punctual features are often overestimated or even missed, particularly in embedding-based methods. DDAD demonstrates qualitatively enhanced reconstruction capability, however, it struggles with patches that exhibit large background variations. Although indicating solid segmentation performance, PRO-scores do not consider FP assignments, thus overestimations of defective regions are not directly penalized.
ASM consists of numerous patches containing imperfect masked image borders, due to varying imaging conditions and subsequent influence on segmentation performance of applied deep learning-based models. Additionally, this modality is characterized by large normal feature variations caused by the underlying imaging procedure, which relies on direct reflection.
All approaches exhibited misclassifications in these regions, often leading to TP assignments due to coincidences of real defective areas within the same patch. As a result, despite achieving low performance among all modalities, overestimations of image-level metrics must be considered. In addition to small punctual defects, it was particularly difficult to detect pronounced squeegee strokes among the methods evaluated. However, the anomaly maps (Fig. 5) indicate that the student-teacher based approach Efficient-AD achieved the best qualitative segmentation performance. Again, the localiced patch descriptor seems preferable for suppressing false positives.
Above conducted experiments revealed that the defect detection performance of leveraged unsupervised approaches is strongly dependent on dataset properties, determined by its underlying imaging modalities, preprocessing and defect appearance.
Showing distinct defective feature characteristics, LSM-2 achieved best overall performance, resulting in industrial applicable FPR of Efficient-AD and U-Flow, altough struggling with weakly contrasted pinholes. Compared to the supervised approach in section 4.1, FPRs are mostly significantly higher (up to factor 10 in ASM), limiting industrial applicability, particularly in the case of ASM.
Despite DDAD’s qualitatively good reconstruction capability, modelling permitted pattern variations and patterns in border regions was often not feasible (Fig. 6). In addition, DDAD exhibited reconstruction instabilities resulting in noise that increased with higher domain adaptation weightings.
5.3.2 Limitations
The evaluations of these approaches should be considered as estimations based on the chosen methods parameters, preprocessing techniques, and training settings, including random seeds. Extensive hyperparameter tuning and advanced preprocessing strategies are left for future research.
6 Conclusions and Outlook
This publication introduces a novel surface defect detection dataset that exhibits a “real-world” industrial use case. The investigated screen-printed samples were captured using three different optical modalities. Typical for industrial inspection, the dataset reflects imperfect imaging conditions caused by imaging modalities, data preprocessing, and permitted sample variance, thereby presenting challenges for defect detection algorithms. In addition to extending common unsupervised settings, the dataset includes large-scale supervised training data, comprising tens of thousands of fault-free patches. Based on a review of recent literature, the proposed dataset can be assumed to be the largest publicly available, enabling research on both unsupervised and supervised approaches by using both synthetic and real defects.
It was observed that SOTA unsupervised methods, which achieve remarkable performance on established benchmark datasets such as MVTec [8], struggled in these specific industrial scenarios. The occurence of small and weakly contrasted defects, combined with high permitted design variability, significantly limited defect detection performance, especially at the pixel-levels. Among the evaluated methods, the most promising approaches are based on defect synthetization at both the feature and image-levels and knowledge distillation using patch descriptor networks, achieving image-level MCCs of up to 0.89 in LSM-1 and 0.96 in LSM-2. However, despite robust performance of Efficient-AD with industry applicable low FPRs in LSM-1 and LSM-2, weakly contrasted defects remained challenging. While certain unsupervised approaches have shown promising advances in image-level detection performance for specific imaging modalities (e.g., LSM-2), future research should focus enhancing segmentation performance under industrial imaging conditions.
Incorporating available weakly labeled defective samples within a mixed supervised training procedure is able to improve image-level defect detection performance by large margins. The experiments described in section 5.2 indicate that the initial comprehensive feature distribution, by means of several tens of thousands of synthesized punctual defects, is complemented and made more diverse, even with small amounts of available real defects. The injection of small fractions (9 samples in ASM) of previously unseen defect classes (area defects) substantially improved model generalization. The defect detection performance of small and weakly contrasted point defects could be significantly improved during mixed training (MCC $=$ 0.96 in LSM-1), compared to training on solely real defects (MCC $=$ 0.91).
These findings suggest a valuable capability for industrial applications, as emerging defects, both seen and previously unseen, can be efficiently integrated into subsequent scalable training, enhancing overall feature diversity. Thus, initial training on synthetic defects can be incrementally improved, meeting process requirements such as a low FPR and high recall.
Furthermore, purely synthetic training achieved near-optimal (MCCs up to 0.99) defect detection performance in modalities that exhibit distinct feature characteristics, such as pinholes in LSM-2. Purely synthetic training outperformed SOTA unsupervised approaches in two out of three modalities (LSM-2 and ASM).
Despite promising investigations of generalization capabilities with a limited number of available defective samples, further experiments on more diverse defect classes with larger test sets are necessary to gain deeper insights.
Additionally, permitted normal data variations are learned through large amounts (50 0005000050\,00050 000100 000100000100\,000100 000 of augmented fault-free patches, even with noisy labels. Noisy labels, due to e.g. process contamination during the image acquisition, were estimated to account for 2 – 5 % in LSM-1. These corrupted labels did not show a significant impact during training. Thus, the screening effort of fault-free data in the proposed unsupervised training datasets could be considered more elaborate than the presented supervised data preprocessing workflow in [35], minimizing manual labeling effort to known defective sample regions. Thus, further research should focus on efficient and industrial applicable data preparation and preprocessing workflows within unsupervised and supervised approaches and the influence of different proportions of noisy labels.
As a direct implication of above findings, research in self-supervised approaches that utilize these representative normal data distributions should be emphasized to enhance industry-applicable model robustness at low labeling costs.
Recent supervised anomaly detection methods [57, 76, 34] align with our findings, demonstrating improved detection performance through the integration of limited labeled data in the mixed training process. Those results indicate that incorporating both synthetic and real defects during training is a promising strategy in industrial defect detection.
6.1 Future Applications
The proposed dataset is designed to advance future research on both unsupervised and supervised approaches suitable for challenging industrial anomaly detection use cases. Leveraging the proposed dataset enables further exploration in the following directions:
  • Investigation of the impact of increased normal feature variance on the robustness of unsupervised approaches by utilizing larger sets of fault-free training samples than those used in the proposed experiments. Additionally, an examination of the effect of real noisy labels by applying the “unscreened” automatically extracted fault-free training data.
  • Conduction of research on self-supervised learning approaches [29] utilizing the synthetic or fault-free large scale training sets. Therefore, feature extractors of unsupervised approaches, such as described in [33, 41], could be pretrained to learn more discriminative features of the target domain.
  • Encouraging research on methods that integrate both synthetic and real defects during training, as demonstrated by the proposed mixed supervised training and recent supervised anomaly detection methods.
[…]
Acknowledgements The research work was performed within the COMET-project: Deep on-line learning for highly adaptable polymer surface inspection systems (project-no.: 879785) at the Polymer Competence Center Leoben GmbH (PCCL, Austria) within the framework of the COMET-program of the Federal Ministry for Climate Action, Environment, Energy, Mobility, Innovation and Technology and the Federal Ministry for Digital and Economic Affairs and with contributions by Burg Design GmbH. The PCCL is funded by the Austrian Government and the State Governments of Styria, Lower Austria and Upper Austria.
[…]
Funding Open access publishing enabled by Montanuniversität Leoben agreement with Springer Nature (Austria KEMÖ: Springer Transformative Agreement). This work was supported within the COMET-project (project-no.: 879785).
[…]
Data availability The dataset supporting this research is publicly available at Zenodo https://doi.org/10.5281/zenodo.14911043.
[…]
Author contribution Conceptualization: PJK and DPG; Methodology: PJK; Formal analysis and investigation: PJK; Writing - original draft preparation: PJK; Writing - review and editing: PJK and DPG; Funding acquisition: DPG; Supervision: DPG.
Declarations
[…]
Competing interests The authors have no competing interests to declare that are relevant to the content of this article.
Appendix A
Table 7: Overview of the available data splits in the proposed ISP-AD dataset. The dataset is divided into unsupervised training data, containing screened fault-free patches, and supervised training data, which includes "noisy" fault-free patches along with additional synthetic and real defective data. Test splits remain the same for both settings, mimicking the imbalanced data distributions typical of industrial inspection. Additional unaugmented fault-free data available is not included in this listing.
Train Test
Unsupervised Supervised Super- / Unsupervised
Modality Good Good #Synthetic #Area #Points Good #Area #Points
LSM-1 3678 100 669100669100\,669100 669 99 3319933199\,33199 331 120 52 1470 49 46
LSM-2 401 50 0365003650\,03650 036 49 9644996449\,96449 964 62 66 382 63 33
ASM 500 103 622103622103\,622103 622 96 3699636996\,36996 369 37 109 1916 33 41
Table 8: Augmentation settings of generated supervised training set including random affine and illumination transformations.
Affine Transformations Illumination Tranformations
Modality Rotation Scale Shear Vertical Flip Horizontal Flip Brightness Contrast
LSM-1 45.0 (0.90,1.10) 10.0 yes yes (0.75,1.25) (0.75,1.25)
LSM-2 45.0 (0.90,1.10) 20.0 yes yes (0.75,1.25) (0.75,1.25)
ASM 10.0 (1.0,1.2) 10.0 yes yes no no
Table 7 illustrates the training and test data splits of the proposed ISP-AD dataset. The augmentation settings applied to the supervised training data are shown in Table 8. In addition to the presented data splits, large-scale fault-free training data (10 $0 0$ 20 0002000020\,00020 000 patches) without augmentation is made available for all three modalities: LSM-1, LSM-2, and ASM. As outlined in Section 3.2, the extracted patches can be considered noisy. The naming convention of image-level labels within the .hdf5 files follows:
  • Stream including both fault-free and synthetic defects: syn_stream
  • Ground truth masks: ground_truth
To reproduce the experiments presented in Section 5.2, randomly sampled fractions of the mixed, area, and points defect groups will be made available separately.
Figure 5: Visualization of the pixel-level detection performance (anomaly map overlay) for the investigated unsupervised methods (Efficient-AD [6], U-Flow [64], GLASS [14], and DDAD [48]) on ASM. Efficient-AD demonstrated the most accurate detection of the depicted punctual defect. The reconstruction-based approach DDAD produced a high number of false positive detections in transmission areas of direct reflection and masked border regions.
Figure 6: Illustration of the reconstruction capabilities of DDAD [48] on a patterned surface of LSM-1. The denoising diffusion process enabled a qualitatively good reconstruction, effectively omitting the small fiber in the raw image. However, fine-grained details, such as contrast variations in the grid regions, remained challenging, resulting in high anomaly scores across broad areas.